john likes attending cinema and theatre the marginal utility he gets from going to the cinema is 20 and the marginal utility he receive from going to the theatre is 60. what can we conclude from his preferences over cinema and theatre.
Answer #1

As marginal utility for theatre is greater than for cinema, John prefers theatre over cinema. Now, plotting cinema on x-axis and movies on y-axis, slope of the tangent to the indifference curve at the given consumption level = - MUT/MUC = -60/20 = -3.